OpenStudy (anonymous):

What is the value of 10c6

OpenStudy (misty1212):

hi!!

OpenStudy (misty1212):

is this \[_{10}C_6\] ten choose six?

OpenStudy (anonymous):

what are the choices?

OpenStudy (misty1212):

same as \[_{10}C_4\] compute via \[\frac{\overbrace{10\times 8\times 7\times 6}^{\text{four terms}}}{4\times 3\times 2}\] cancel first multiply last

OpenStudy (anonymous):

10C6 = 10! / (6! * 4!) = 10 * 9 * 8 * 7 / 4! = 5040/24 = 210.

OpenStudy (misty1212):

oops \[\frac{\overbrace{10\times 9\times 8\times 7}^{\text{four terms}}}{4\times 3\times 2}\]

OpenStudy (anonymous):

Lol yeah now thats right :) @misty1212